# On Memorization of Large Language Models in Logical Reasoning

### v1

@@ 4.1 Reasoning capabilities of Direct FT-ed model
Figure 7: Test accuracy improvement on $N$-people problems for LLMs fine-tuned on $M$-people problems, compared to the unfine-tuned model, under 0-shot direct prompting. Most grid values are above 0, indicating transferability and enhanced reasoning abilities across unseen tasks. Results for more epochs are in § C.2.
@@
Fine-tuned model generalizes across different difficulty levels. We evaluate LLMs’ transferability by fine-tuning on $M$-people puzzles and testing on $N$-people puzzles. When $M\neq N$, the testing is out-of-distribution compared to training and solving it requires reasoning. The $N\times M$ test accuracy improvement grid (compared to the un-FTed model) in Fig. 7 shows: (1) Training on any $M$-people puzzle generally improves test accuracy on any $N$-people puzzles, suggesting that the model learns general task-solving rules after FT (to reason and solve both easier and harder unseen puzzles). (2) More training epochs (e.g., 50 vs. 5) improve results, especially for Llama3-8B. (3) Accuracy gains are larger for $N\leq 6$ puzzles, though improvements on harder tasks remain possible.
@@
Fine-tuning with 10k 8-people puzzles. Given the significant performance improvement from fine-tuning, a natural question arises: can brute-force fine-tuning on a very large number of puzzles eventually solve the K&K puzzles, by observing/memorizing a variety of combinations of persons’ claims and their corresponding answers? We Direct FT GPT4o-mini on $1k$/$10k$ of the most challenging 8-people puzzles for 5 epochs. Fig. 8 shows that (1) $10k$-FT significantly outperforms $1k$-FT across all tasks, reaching $\sim 90\%$ test accuracy on moderately difficult 4/5-people puzzles. (2) CoT FT is generally more effective than Direct FT with $10k$ samples, likely due to the guidance provided by reasoning steps. (3) An exception is the 2-people task, where the training and testing distribution gap causes the CoT FTed model to occasionally get stuck in a loop of listing assumptions and contradictions, resulting in long, repetitive responses without reaching a conclusion [^2] We observe similar accuracy drop on 2-people task for Llama3-8B (see Fig. 22) when it is Direct FTed for overly long epochs. We provide more examples and discussions in § C.2.2.. (4) Direct FT with $10k$ puzzles achieves surprisingly high test accuracy on all tasks, e.g., 52% on 8-people tasks, where the un-FTed model scores near 0. Notably, the models do not see reasoning steps during training and rely solely on memorizing answers. We also observe high transferability for $10k$ Direct FTed Llama3-8B in Fig. 22, e.g., $87\%$ test accuracy on 3-people puzzles.
@@ 4.2 Probing Direct FTed models
To investigate whether Direct FTed models develop internal understanding of the skills necessary to solve K&K puzzles when learning only from the answers, we use probing techniques (Adi et al., 2017, Conneau et al., 2018, Hewitt & Liang, 2019, Ye et al., 2024) to analyze their internal representations. Specifically, we study whether a Direct FTed model’s intermediate outputs provide evidence that it can distinguish between correct and incorrect statements for a given K&K puzzle, which is essential for solving the puzzle via reasoning. For a given model, we extract intermediate outputs from all transformer blocks for 200 correct and 200 incorrect statements, then check whether these outputs form distinct clusters by measuring the training accuracy of a logistic regression model fit on them (see § B.2.3 for details). For each $N$-people K&K puzzle, we report the per-layer probing accuracy averaged across seven Direct FTed models, each Direct FTed on an $M\in\{2,3,\ldots,8\}$-people task.
Figure 9: Probing accuracy of K&K puzzles with different number of people in testing puzzles across different layers of the Llama3-8B transformer model. Results for un-FTed models are shown in Fig. 31 in § C.
Fig. 9 shows (1) a clear trend of higher probing accuracy in deeper layers, peaking at around the 14th/15th layer. The near-perfect peak accuracy suggests that the model’s internal representations have a clear distinction between true/false statements about a given puzzle. (2) The probing accuracy is much higher than the un-FTed model (Fig. 31 in § C), suggesting that such representations are learned from the question-answer pairs during Direct FT. (3) Puzzles with more people seem to demand more internal computation, as evidenced by the point where probing accuracy surpasses $85\%$ shifting to later transformer blocks.
@@ 4.3 Direct FT with Wrong Answers
Figure 10: Test accuracy improves on $N$-people puzzles for Llama3-8B fine-tuned on $M$-people puzzles with completely wrong answers, compared to the unfine-tuned model. This evaluation uses 1-shot direct prompting (see Fig. 27 for results under different prompting setups).
Note that in this case the improved test accuracy could not have come from memorization because 100% of the training examples are incorrectly labeled. However, since in each wrong answer of a $N$-people puzzle, there are still $N-\tilde{N}$ correct role assignments where the random $\tilde{N}\geq 1$. The model might have learned to reason from those partially correct role assignments in the wrong answer.
@@
However, as shown in Fig. 11, when applied to more capable GPT4o-mini models, Direct FT on 5-people puzzles where 100% training examples have corrupted answers does not lead to improvement. Moreover, the negative effects transfer to other tasks, notably easier ones (2/3/4-people). Nevertheless, as the percentage of corrupt-answer training examples reduces ($\leq 50\%$), the model could gain improved reasoning capabilities that generalize across different $N$-people tasks. We provide GPT4o-mini results under more epochs in Fig. 30 and Llama3-8B results for partially wrong answer FT in Figs. 28 and 29.
5 LLMs Learn to Reason by Fine-tuning with CoTs
@@
Model learns to reason on CoT when model capacity is large enough. As shown in Fig. 5, (1) training with reasoning steps as guidance improves test accuracy ($y$-axis) on unseen puzzles. (2) However, Llama3-8B struggles with CoT FT, likely due to its limited capacity to effectively learn CoT skills with $\leq$1K training samples. (3) Similar to Direct FT results in § 4, in CoT FT, memorization of training data is higher than test data (Fig. 5), yet test accuracy improves despite that the memorization score increases over training (Fig. 18), and the fine-tuned models show positive transferability to easier/harder tasks (Fig. 7). (4) Though models can generalize surprisingly well under Direct FT, CoT FT could lead to much higher test accuracy, especially with a larger training set (Fig. 8).
@@
6 Distinguishing Memorization from Reasoning
The findings above show that models’ reasoning capabilities continue to improve as they memorize more training examples. In other words, the models use both memorization and reasoning to solve the puzzles. How do models decide which example to memorize or reason about? We can use our framework to study this question by extending the memorization score to a per-example metric.
Specifically, consider measuring Eq. 1 on a 1-point dataset $\mathcal{D}=\{x\}$. We skip the examples where $\mathsf{Acc}(f;\{x\})=0$ as the consistency ratio $\mathsf{CR}(\{x\})$ is NaN in this case. Then $\mathsf{LiMem}(f;\{x\})\in\{0,1\}$ is a binary indicator: 0 indicates $x$ is consistently solved after local perturbation; 1 indicates otherwise. We would like to see if there is a clear rule that can separate the two types of puzzles.
Setup. We collect training samples $\{x_{i}\}$ on which the targeted LLM makes correct predictions, and assign a binary categorical label as either “consistently solved” (i.e., solved by reasoning) puzzle or “not consistently solved” (i.e., solved by memorization) puzzle. We split this dataset into random disjoint 80%/20% training/test sets, and train a simple logistic regression model to solve this binary classification problem, in order to study the question: Is there a simple indicator that determines whether a model would solve a given puzzle by reasoning or memorization?
@@ 6 Distinguishing Memorization from Reasoning
Puzzle-based indicators. We consider the following features: (1) TF-IDF; (2) Bag-of-Words; (3) Word Length; (4) Character Length; (5) concatenation of all. Each feature can be extracted from one of the following fields: (1) question; (2) synthetic CoT reasoning steps; (3) model response [^3] Strictly speaking this is a model-based indicator feature.; (4) concatenation of the above fields. The training and test performance (measured with AUC as the dataset can be unbalanced) are shown in Fig. 13. We observe a test AUC of 0.629/0.787 for Direct/CoT FT-ed GPT4o-mini, and 0.627 for Direct FT-ed Llama3-8B. This indicates that the puzzle-based indicators could be informative, though not perfect, at determining which examples are reasoned vs. memorized.
Model-based indicators. Here we study model-based indicators to test whether the internal activations of the fine-tuned model are informative for this categorization. Since we do not have access to the model internals of GPT4o-mini, we conduct the experiment on Llama3-8B. Specifically, we feed each puzzle question to the FT-ed model, collect the average embedding at each layer as features, and then train a linear classifier based on the features of each layer. The test AUCs are shown in Fig. 14, where we compare the results based on a not-FTed base Llama3-8B model, to see how much of the feature expressiveness comes from the fine-tuning process. We observe that (1) the features from the bottom layers are too low level to classify memorization vs. reasoning, but the test AUCs improve at higher layers. (2) The features from the FTed model are consistently more informative than the un-FTed model, suggesting that the model’s decision regarding memorization vs. reasoning on specific samples likely stems from the fine-tuning process. (3) The best model embedding-based indicator provides stronger signals than the puzzle-based indicator (Fig. 13 left) for Llama3-8B, with 0.70 and 0.627 test AUC on 3-people puzzles, respectively.
@@ 7 Related Work
Memorization in LLMs. Prior work has explored training data memorization in LLMs, primarily in the contexts of privacy and copyright concerns (Carlini et al., 2021, Lukas et al., 2023, He et al., 2024), focusing on how LLMs may reproduce text near-verbatim to their training data (Lee et al., 2022, Carlini et al., 2023, Biderman et al., 2024). In contrast, we examine memorization in the reasoning context, and focus on analyzing whether LLMs can accurately solve problems encountered during training but struggle to solve slightly perturbed variants. This allows us to better investigate the extent to which LLMs truly understand and generalize the underlying principles of the reasoning problems they have been trained on, as opposed to merely memorizing the text.
Recent research discusses signs of LLMs memorization in reasoning tasks by evaluating them on counterfactual reasoning tasks. These counterfactual tasks demand similar abstract reasoning skills as the original tasks but are less common in the training data. For instance, tasks such as reversing a sequence of words (McCoy et al., 2024) show better performance on high-probability sequences than on low-probability sequences; shifting each letter by $n$ places in the alphabet (Rot-$n$) (Prabhakar et al., 2024, McCoy et al., 2024) demonstrates higher performance when $n=13$ than for other values, likely because “Rot-13” is commonly used in online forums.  Wu et al. (2024b) presents 11 counterfactual tasks (e.g., 1-indexing in Python, base-9 arithmetic) that show significant performance declines. Jiang et al. (2024) changes some tokens in the reasoning task descriptions which leads to significant performance drops, suggesting that models might depend on recognizing superficial patterns with strong token bias. Moreover, Razeghi et al. (2022) finds a strong correlation between the accuracy for a number on numerical reasoning tasks and its frequency in pretraining for GPT-J/GPT-Neo. In our study, we formally define a memorization score to quantify performance variance under task perturbations, covering both counterfactual alterations (e.g., switching the roles of knights and knaves) and standard perturbations on language level and problem structure level.
Detecting benchmark contamination. Recent work has shown that LLMs’ performance drastically declines when faced with altered versions of popular reasoning benchmarks, suggesting potential contamination/memorization of these benchmarks. The benchmark variants include diverse forms such as altered multiple-choice questions formats (Wang et al., 2024, Zong et al., 2024, Gupta et al., 2024, Zhou et al., 2024, Robinson & Wingate, 2023), rephrased or translated problems (Xu et al., 2024, Yang et al., 2023, Yao et al., 2024), shuffled example orderings (Oren et al., 2024), human-curated problems of comparable difficulty (Zhang et al., 2024), functional variants generating random instantiations (Srivastava et al., 2024, Mirzadeh et al., 2024), and problems beyond specific date cutoffs (Roberts et al., 2023, Jain et al., 2024). Previous work either focus on surface level language perturbations or require extensive expert-level annotations for math level variations. In contrast, our benchmark support automatic problem-level perturbation, solution and reasoning procedure generation, and easily scale to different difficult levels and dataset sizes without extra human efforts.
Logical reasoning benchmarks. To evaluate logical reasoning capabilities in LLMs, synthetic benchmarks have been developed to enable scalable generation of samples with varying configurations and difficulty levels (Clark et al., 2020, Giadikiaroglou et al., 2024, Parmar et al., 2024). For instance, DyVal (Zhu et al., 2024) uses directed acyclic graphs to dynamically generate samples on reasoning tasks including deductive, Boolean, and abductive reasoning. Chen et al. (2024) focus on propositional logical problems involving definite clauses, and synthetically generate variations with different premise orders, such as forward, backward, and shuffled. Dziri et al. (2024) explore the limitations of LLMs in tasks requiring compositional reasoning, including multiplication, logic grid puzzles, and dynamic programming problems. ZebraLogic (Lin et al., 2024) is an extended benchmark that systematically tests logical reasoning capabilities. BoardgameQA (Kazemi et al., 2024) presents a question-answering dataset characterized by contradictory facts and rules in the questions. TruthQuest (Mondorf & Plank, 2024) is the most similar task to our work, which provides evaluation samples based on K&K-type of puzzles involving 3-6 person. Our work provides more comprehensive dynamic set of K&K puzzles that support automatic generation of perturbations, solutions and detailed reasoning steps. Moreover, based on this benchmark, we define and measure memorization in reasoning tasks, revealing intricate interplay between memorization and reasoning in LLMs.
Improving reasoning via fine-tuning. Prior work has explored fine-tuning LLMs on synthetic reasoning data to enhance their performance on reasoning. DyVal (Zhu et al., 2024) shows that fine-tuning Llama2-13B-chat on their synthetic reasoning benchmark improves its performance on other popular reasoning benchmarks. BoardgameQA (Kazemi et al., 2024) find that fine-tuning BERT-large and T5-XXL on their training dataset with synthetic proofs outperforms few-shot CoT prompting using PaLM. Ye et al. (2024) pretrain GPT2 from scratch on synthetic math problems, synthetic CoT steps and solutions and show that model can solve problems from the same distribution and generalize to out-of-distribution (OOD) problems. However, Dziri et al. (2024) show that while GPT-3 fine-tuned on their compositional reasoning tasks with/without reasoning steps can solve in-distribution (ID) problems, it fails to generalize to OOD tasks with increased problem sizes. Besides using synthetic CoTs, there are work using model-generated CoTs to enhance the models’ reasoning capabilities (Chung et al., 2024). STaR (Zelikman et al., 2022) uses model self-generated CoTs on correctly solved samples to iteratively fine-tune itself as a self-taught reasoner. A number of work (Puerto et al., 2024, Kim et al., 2023, Ho et al., 2023, Hsieh et al., 2023) leverage CoTs generated from teacher models to train smaller student models. Additonally, some recent efforts have focused on leveraging intermediate reasoning steps in CoT more implicitly. For instance, Deng et al. (2023) distill intermediate reasoning tokens into the network layers by representing reasoning steps as vectors and using them as targets; Deng et al. (2024) distill CoT by gradually removing the intermediate steps and fine-tuning the model to internalize these steps, predicting the answers based on partial CoT. Both studies show that full CoT fine-tuning may not be necessary for the model to achieve strong reasoning performance.
8 Conclusion and Future Work

### v2

@@ 2.2 Knights and Knaves Reasoning Benchmark
Figure 6: Test accuracy improvement on $N$-people problems for LLMs fine-tuned on $M$-people problems, compared to the unfine-tuned model, under 0-shot direct prompting. Most grid values are above 0, indicating transferability and enhanced reasoning abilities across unseen tasks. Results for more epochs are in § E.3.
3 Quantifying Memorization in Reasoning
Fine-tuning setup. We fine-tune the models for each $N$-people task separately, with $n_{\text{train}}=1,000$ for $3\leq N\leq 8$, and $n_{\text{train}}=200$ for $2$-people task due to limited number of combinations. We take Llama3-8B and GPT4o-mini and run supervised fine-tuning (SFT) on a set of K&K training puzzles disjoint from the test set. We consider two fine-tuning paradigms: (1) Fine-tuning on detailed CoT steps (CoT FT): during SFT, the model observes the concatenation of the question, synthetic CoT steps, and the answer for each puzzle; the loss is computed on the CoT steps and the answer part. (2) Fine-tuning on the answers (Direct FT) where the model observes the question-answer pair for each puzzle, and the loss is only computed on the answer part. Examples of CoT FT/Direct FT training instances are provided in § D.2.2. We fine-tune Llama3-8B for $50$ epochs [^3] We fine-tune Llama3-8B for max 100 epochs in Fig. 26 and find that it typically converges at 50 epochs. and GPT4o-mini for $5$ epochs (due to budget constraints) via the OpenAI fine-tune API (details in § D.2). During the evaluation, we follow the same prompting paradigm as FT paradigm, i.e., direct/CoT prompting for direct/CoT-FTed model, which is shown effective in § E.3.
LLMs interpolate K&K training puzzles. In Fig. 5, we present the training accuracy of models trained on each task on the $x$-axis (each dot represents a training epoch). We find that models exhibit high training accuracy in tasks such as $3/5$-people puzzles. The higher capacity model GPT4o-mini nearly achieves interpolation ($\mathsf{Acc}(f;\mathsf{Tr})\approx 100\%$) using both Direct FT and CoT FT.
4 Learn to Reason by Fine-tuning With Answers Only
Figure 7: Inconsistency ratio (y-axis) on correctly solved training puzzles of fine-tuned Llama3-8B decreases over epochs (x-axis), even as the proportion of memorized training puzzles increases, as indicated by the larger $\mathsf{LiMem}(f;\mathsf{Tr})$ values (color bar).
Figure 8: Probing accuracy of K&K puzzles with different number of people in testing puzzles across different layers of the Llama3-8B transformer model. Results for un-FTed models are shown in Fig. 38 in § E.
§ 3 shows that fine-tuned models exhibit memorization when solving K&K reasoning tasks. Does it mean that those models do not have reasoning capabilities at all? In this section, we show that LLMs can do both, and the reasoning capability consistently improves as the memorization level increases when the models are fine-tuned on K&K puzzles.
We focus on analyzing Direct FT in this section and discuss CoT FT in § 5. For humans, solving K&K tasks without understanding the underlying logic is difficult. However, after observing the step-by-step reasoning steps, people can understand the procedure and solve the puzzles more easily. Similarly, compared to CoT FT, learning from only answers (Direct FT) without detailed reasoning steps is intuitively more challenging for LLMs, as the models need to come up with the reasoning procedures on their own. Therefore, the models might be more likely to rely on memorization in this case. Surprisingly, from Fig. 5, we did not observe Direct FTed GPT4o-mini models exhibiting consistently higher memorization score than CoT FTed ones. It turns out that models can learn to reason K&K puzzles well directly from observing only question-answer pairs, as we will show in § 4.1. To better understand what the model learns through Direct FT, we conduct a probing analysis on model internals in § 4.3 and an ablation study with incorrect answers fine-tuning in § 4.3.
4.1 Reasoning capabilities of Direct FT-ed model
@@
Fine-tuned model generalizes across different difficulty levels. We evaluate LLMs’ transferability by fine-tuning on $M$-people puzzles and testing on $N$-people puzzles. When $M\neq N$, the testing is out-of-distribution compared to training and solving it requires reasoning. The $N\times M$ test accuracy improvement grid (compared to the un-FTed model) in Fig. 6 shows: (1) Training on any $M$-people puzzle generally improves test accuracy on any $N$-people puzzles, suggesting that the model learns general task-solving rules after FT (to reason and solve both easier and harder unseen puzzles). (2) More training epochs (e.g., 50 vs. 5) improve generalization, especially for Llama3-8B. (3) Accuracy gains are larger for $N\leq 6$ puzzles, though improvements on harder tasks remain possible.
Inconsistency ratio decreases despite increased memorization. As shown in Fig. 7, the inconsistency ratio ($y$-axis) of fine-tuned LLMs on correctly solved training puzzles decreases over epochs, even as the memorization score $\mathsf{LiMem}(f;\mathsf{Tr})$ increases, indicating a higher proportion of memorized training puzzles (Eq. 1). This reduction in inconsistency suggests a potential improvement in the model’s generalization ability, aligning with its enhanced transferability observed in Fig. 6. The memorization score $\mathsf{LiMem}(f;\mathsf{Tr})$ under role-flipping is significantly higher than other perturbation, possibly due to an internal bias that knights are truthful. See Fig. 22 for results on GPT4o-mini.
@@
Fine-tuning with 10k 8-people puzzles. Given the significant performance improvement from fine-tuning, a natural question arises: can brute-force fine-tuning on a very large number of puzzles eventually solve the K&K puzzles, by observing/memorizing a variety of combinations of persons’ claims and their corresponding answers? We Direct FT GPT4o-mini on $1k$/$10k$ of the most challenging 8-people puzzles for 5 epochs. Fig. 9 shows that (1) $10k$-FT significantly outperforms $1k$-FT across all tasks, reaching $\sim 90\%$ test accuracy on moderately difficult 4/5-people puzzles. (2) CoT FT is generally more effective than Direct FT with $10k$ samples, likely due to the guidance provided by reasoning steps. (3) An exception is the 2-people task, where the training and testing distribution gap causes the CoT FTed model to occasionally get stuck in a loop of listing assumptions and contradictions, resulting in long, repetitive responses without reaching a conclusion [^4] We observe similar accuracy drop on 2-people task for Llama3-8B (see Fig. 29) when it is Direct FTed for overly long epochs. We provide more examples and discussions in § E.3.2.. (4) Direct FT with $10k$ puzzles achieves surprisingly high test accuracy on all tasks, e.g., 52% on 8-people tasks, where the un-FTed model scores near 0. Notably, the models do not see reasoning steps during training and rely solely on memorizing answers. We also observe high transferability for $10k$ Direct FTed Llama3-8B in Fig. 29, e.g., $87\%$ test accuracy on 3-people puzzles.
@@ 4.2 Probing Direct FTed models
To investigate whether Direct FTed models develop internal understanding of the skills necessary to solve K&K puzzles when learning only from the answers, we use probing techniques (Adi et al., 2017; Conneau et al., 2018; Hewitt & Liang, 2019; Ye et al., 2024) to analyze their internal representations. Specifically, we study whether a Direct FTed model’s intermediate outputs provide evidence that it can distinguish between correct and incorrect statements for a given K&K puzzle, which is essential for solving the puzzle via reasoning. For a given model, we extract intermediate outputs from all transformer blocks for 200 correct and 200 incorrect statements, then check whether these outputs form distinct clusters by measuring the training accuracy of a logistic regression model fit on them (see § D.2.3 for details). For each $N$-people K&K puzzle, we report the per-layer probing accuracy averaged across seven Direct FTed models, each FTed on an $M\in\{2,\ldots,8\}$-people task.
Fig. 8 shows (1) a clear trend of higher probing accuracy in deeper layers, peaking at around the 14th/15th layer. The near-perfect peak accuracy suggests that the model’s internal representations have a clear distinction between true/false statements about a given puzzle. (2) The probing accuracy is much higher than the un-FTed model (Fig. 38 in § E), suggesting that such representations are learned from the question-answer pairs during Direct FT. (3) Puzzles with more people seem to demand more internal computation, as evidenced by the point where probing accuracy surpasses $85\%$ shifting to later transformer blocks.
@@ 4.3 Direct FT with Wrong Answers
Figure 10: Test accuracy improves on $N$-people puzzles for Llama3-8B Direct FTed on $M$-people puzzles with completely wrong answers, compared to the unfine-tuned model. This evaluation uses 1-shot direct prompting (see Fig. 34 for results under different prompting setups).
Note that in this case the improved test accuracy could not have come from pure memorization because 100% of the training examples are incorrectly labeled. However, since in each wrong answer of a $N$-people puzzle, there are still $N-\tilde{N}$ correct role assignments where the random $\tilde{N}\geq 1$. The model might have learned to reason from those partially correct role assignments in the wrong answer.
@@
However, as shown in Fig. 11, when applied to more capable model GPT4o-mini, Direct FT on 5-people puzzles where 100% training examples have corrupted answers does not lead to improvement. Moreover, the negative effects transfer to other tasks, notably easier ones (2/3/4-people). Nevertheless, as the percentage of corrupt-answer training examples reduces ($\leq 50\%$), the model could gain improved reasoning capabilities that generalize across different $N$-people tasks. We provide GPT4o-mini results under more epochs in Fig. 37 and Llama3-8B results for partially wrong answer FT in Figs. 35 and 36.
5 Learn to Reason by Fine-tuning with CoTs
@@
Model learns to reason on CoT when model capacity is large enough. As shown in Fig. 5, (1) training with reasoning steps as guidance improves test accuracy ($y$-axis) on unseen puzzles. (2) However, Llama3-8B struggles with CoT FT, likely due to its limited capacity to effectively learn CoT skills with $\leq$1K training samples. (3) Similar to Direct FT results in § 4, in CoT FT, memorization of training data is higher than test data (Fig. 5), yet inconsistency ratio decreases despite that overall memorization score increases over training (Fig. 22), and the fine-tuned models show positive transferability to easier/harder tasks (Fig. 6). (4) Though LLMs can generalize surprisingly well under Direct FT, CoT FT could lead to much higher test accuracy, especially with a larger training set (Fig. 9).
@@
6 Distinguish Memorization from Reasoning
The findings above show that models’ reasoning capabilities continue to improve as they memorize more training examples. In other words, the models use both memorization and reasoning to solve the puzzles. But how do they decide which examples to memorize or reason about? We can explore this by extending our memorization score to a per-example metric. Specifically, consider measuring Eq. 1 on a 1-point dataset $\mathcal{D}=\{x\}$. We skip the examples where $\mathsf{Acc}(f;\{x\})=0$ as the consistency ratio $\mathsf{CR}(\{x\})$ is NaN in this case. Here, $\mathsf{LiMem}(f;\{x\})\in{0,1}$ acts as a binary indicator: 0 means $x$ remains solvable after local perturbation, while 1 means it does not. Our goal is to determine if a clear rule separates these two types of puzzles.
Setup. We collect training samples where the targeted LLM predicts correctly, and label each as either “consistently solved” (reasoning) or “not consistently solved” (memorization). We then train a logistic regression classifier using an 8:2 train/test split to distinguish between puzzles the model solves through reasoning versus memorization.
@@ 6 Distinguish Memorization from Reasoning
Puzzle-based indicators. We consider the following features: (1) TF-IDF; (2) Bag-of-Words; (3) Word Length; (4) Character Length; (5) concatenation of all. Each feature can be extracted from one of the following fields: (1) question; (2) synthetic CoT reasoning steps; (3) model response [^5] Strictly speaking, this is a model-based indicator feature.; (4) concatenation of the above fields. The train and test performance (measured with AUC as the dataset can be unbalanced) are shown in Fig. 13. We observe a test AUC of 0.629/0.787 for Direct/CoT FT-ed GPT4o-mini, and 0.627 for Direct FT-ed Llama3-8B. This indicates that the puzzle-based indicators could be informative, though not perfect, at determining which examples are reasoned vs. memorized.
Model-based indicators. We also study model-based indicators to test whether the internal activations of the fine-tuned model are informative for this categorization. Since we cannot access model internals of GPT4o-mini, we conduct the experiment on Llama3-8B. Specifically, we feed each puzzle question into the fine-tuned model, extract average embeddings from each layer, and train a linear classifier per layer. Appendix Fig. 40 shows test AUCs, where we also compare the fine-tuned model to its non-fine-tuned counterpart. We observe that (1) Lower-layer features poorly distinguish memorization from reasoning, but higher layers improve. (2) The features from the FTed model are consistently more informative than the un-FTed one, suggesting that the model’s decision regarding memorization vs. reasoning on specific samples likely stems from the fine-tuning process. (3) The best embedding-based indicator (0.70 AUC) outperforms the puzzle-based indicator (Fig. 13 left, 0.627 AUC) on 3-people puzzles.
@@ 7 Related Work
Memorization in LLMs. Previous work on LLM memorization primarily focused on near-verbatim training text regurgitation from the perspective of privacy or copyright concerns (Carlini et al., 2021; Lee et al., 2022; Carlini et al., 2023; Lukas et al., 2023; Biderman et al., 2024; Prashanth et al., 2024). In contrast, we focus on quantifying the memorization behavior of LLMs when solving reasoning tasks, using a metric computed with the help of local perturbation of reasoning puzzles.
Benchmark Contamination and Logical Reasoning Evaluation. Recent research has revealed LLMs’ significant performance decline when faced with altered versions of popular reasoning benchmarks (Oren et al., 2024; Xu et al., 2024; Yang et al., 2023; Yao et al., 2024; Zhang et al., 2024; Srivastava et al., 2024), suggesting potential benchmark contamination. Various synthetic benchmarks have been developed to evaluate LLMs’ logical reasoning capabilities, allowing for dynamic and scalable generation of samples with different configurations and difficulty levels (Clark et al., 2020; Giadikiaroglou et al., 2024; Parmar et al., 2024; Dziri et al., 2024; Lin et al., 2024; Kazemi et al., 2024; Mondorf & Plank, 2024). TruthQuest (Mondorf & Plank, 2024) is the most similar task to our work, which provides K&K-type of 3-6 person puzzles and answers. Our work provides more comprehensive dynamic set of K&K puzzles that support the automatic generation of perturbations, solutions, and detailed reasoning steps. Moreover, we define and measure memorization, and reveal its intricate relation to reasoning. We refer the readers to § B for a more comprehensive discussion of related work (e.g., grokking and adversarial perturbations).
8 Conclusion